OpenStudy (mony01):

Each of the following IVPs corresponds to a harmonic oscillator. Find the solution and specify its amplitude, phase angle and period of motion. x¨ + 9x = 0, x(0) = 1, x˙(0) = 1

OpenStudy (mony01):

@myininaya do you happen to know how to do this?

myininaya (myininaya):

is that a differential equation?

myininaya (myininaya):

just want to make sure i'm seeing right because we were just talking about trig and it seems like we went to diff equations

myininaya (myininaya):

first of all do you know how to solve r^2+9=0

OpenStudy (mony01):

r=3

myininaya (myininaya):

well now exactly r^2-9=0 would have given r=3 or r=-3

myininaya (myininaya):

not exactly* r^2+9=0 would give r=3i or r=-3i

myininaya (myininaya):

so the solution to x''+9x=0 is \[x(t)=A \cos(3t)+B \sin(3t)\]

myininaya (myininaya):

use your conditions to find A and B

myininaya (myininaya):

x(0)=1 x'(0)=1

OpenStudy (mony01):

what do you mean?

myininaya (myininaya):

x(0)=1 means x is 1 when t is 0 x'(0)=1 means x' is 1 when t is 0

myininaya (myininaya):

You will have to find x' give x=Acos(3t)+Bsin(3t)

myininaya (myininaya):

to apply the other condition there

myininaya (myininaya):

you should end up with a system of equations to solve

myininaya (myininaya):

x(0)=1 means x=1 when t=0 so we have \[1=A \cos(3(0))+B \sin(3(0))\] 1=A+B(0) 1=A This should make your life easier when finding B now

myininaya (myininaya):

and it should also show you better how to apply x'(0)=1

myininaya (myininaya):

can you find x' given x=cos(3t)+Bsin(3t) ?

OpenStudy (mony01):

1

myininaya (myininaya):

\[x' \text{ or also called } \frac{dx}{dt}=(\cos(3t))'+(B \sin(3t))'\]

myininaya (myininaya):

to find the derivative of either cos(3t) or sin(3t) you will need chain rule

myininaya (myininaya):

hey @mony01 I'm so sorry I have to leave you

myininaya (myininaya):

@PaxPolaris when you get chance will you help mony

OpenStudy (paxpolaris):

@myininaya Aye Captain my captain! I'll try!

OpenStudy (paxpolaris):

@mony01 ... where are we stuck ?

OpenStudy (mony01):

well i think i did the chain rule right i got -sin(3t)(3)+Bcos(3t)(3)

OpenStudy (paxpolaris):

i feel the need to start over ...by guessing \[x(t)=A \cos \left( 3t+\phi \right)\]

OpenStudy (paxpolaris):

\[x(0)=A \cos(\phi)=1\]\[x'(0)=-3A \sin(\phi)=1\] \[\implies \tan (\phi)=-1/3\]

OpenStudy (mony01):

how did you get -3?

OpenStudy (paxpolaris):

\[x'(t)=A \left[ \cos \left( 3t+\phi \right) \right]'\\=A \left[ -\sin \left( 3t+\phi \right) \right]\cdot3\]

OpenStudy (paxpolaris):

the derivative of cos is negative sin...

OpenStudy (mony01):

oh ok

OpenStudy (paxpolaris):

so phase \(=\phi=-\tan^{-1}(-1/3)=-0.32\)radians

OpenStudy (paxpolaris):

and from that you can derive Amplitude \(=A=\sqrt{10}/3\)

OpenStudy (mony01):

how is tan -1/3 and where did the 3t go from the derivative?

OpenStudy (paxpolaris):

at t=0 we are given the initial conditions x(0)=1 x'(0)=1 because t=0, 3t disappears

OpenStudy (paxpolaris):

And period we already know to be \(2\pi/3\)

OpenStudy (mony01):

how you know tan=-1/3 and where you get 10 and 2pi/3....sorry but im so lost

OpenStudy (paxpolaris):

\[x(0)=A \cos(\phi)=1\]\[x'(0)=-3A \sin(\phi)=1\] do you get these 2 equations?

OpenStudy (mony01):

how did you get those equations?

OpenStudy (paxpolaris):

x(0)=1 and x'(0)=1 ... is given, correct? if \(x(t)=A\cos(3t+\phi)\)... just plug in t=0

OpenStudy (mony01):

ok got that

OpenStudy (paxpolaris):

same thing for x'(0) . ... then divide the second equation from the first to get rid of A\[{-3A \sin(\phi) \over A \cos(\phi)}=\frac11\]

OpenStudy (mony01):

whats the equation for x'(0)?

OpenStudy (mony01):

ok never mind now i see why tan is -1/3

OpenStudy (paxpolaris):

\[x(t)=A \cos(3t+\phi)\\ x'(t)=-3A \sin(3t+\phi)\] \(x'(0)=-3A \sin(\phi)\)...... and \(x'(0)=1\)

OpenStudy (paxpolaris):

ok. so we get the phase angle \(\phi\) ....  and \(cos\ \phi=3/\sqrt{10}\)

OpenStudy (mony01):

why is it sqrt10

OpenStudy (paxpolaris):

its a right triangle (pythagorean theorem) \[\sqrt{(-1)^2+3^2}\]

OpenStudy (mony01):

ok

OpenStudy (paxpolaris):

since \(A\cos(\phi)=1, \implies A=1/\cos(\phi),\implies A=\sqrt{10}/3\)

OpenStudy (mony01):

earlier you said the phase angle was -tan^-1 (-1/3) why?

OpenStudy (paxpolaris):

sorry an extra negative sign. \(\phi\) is the phase angle, and \(\tan(\phi)=-1/3\) .... so \(\phi=...\)

OpenStudy (mony01):

tan^-1 (-1/3)

OpenStudy (paxpolaris):

right so we have the equation: \[x(t)={\sqrt{10}\over3}\cos \left( 3t-0.32.... \right)\] we have the phase angle, and amplitude, ...it fits all the original requirements

OpenStudy (paxpolaris):

what else do we need?

OpenStudy (mony01):

period of motion

OpenStudy (mony01):

for the radians i got -0.32

OpenStudy (paxpolaris):

cos(t) has period 2pi ... correct?

OpenStudy (mony01):

yes

OpenStudy (paxpolaris):

the general function:\[A \cos(\omega\cdot t+\phi)\] has the period: \(2\pi/\omega\)

OpenStudy (mony01):

so the period is 2pi/3

OpenStudy (paxpolaris):

right .... another way to think of it is cos(t) completes 1 period in 2pi cos(3t) completes 3 periods in 2pi

OpenStudy (mony01):

the radians has to be positive right?

OpenStudy (paxpolaris):

idk about that.... guess technically the phase angle can be \(\tan^{-1}(-1/3)+2\pi \cdot n\)

OpenStudy (paxpolaris):

for any integer n

OpenStudy (mony01):

so its always positive?

OpenStudy (paxpolaris):

no i think you can leave it as it is ...

OpenStudy (mony01):

so -.32?

OpenStudy (paxpolaris):

yeah

OpenStudy (mony01):

if it asks to find the simple harmonic motion for the equation, what would it be?

OpenStudy (paxpolaris):

gahhhhhhhhh...........

OpenStudy (paxpolaris):

i've been leading you completely wrong for the last hour. Sorry

OpenStudy (mony01):

what do you mean?

OpenStudy (paxpolaris):

forget everything i said ... go back to what myininaya was doing

OpenStudy (mony01):

i think you were doing it right, i was asking how would i do it if it asked for the harmonic motion

OpenStudy (paxpolaris):

sorry for Panicking ... was checking the answer on wolframalpha ... but initially my equation looked nothing like what they had. it is on correct form

OpenStudy (paxpolaris):

The equation is correct ... one of many ways to write it We got:\[x(t)={\sqrt{10}\over3}\cos \left( 3t-\tan^{-1} \left( \frac13 \right)\right)\] it can also be written using sin instead of cos [BUT the phase angle will be different] \[x(t)={\sqrt{10}\over3}\color{blue}\sin\left( 3t\color{blue}{+\tan^{-1} \left( 3 \right)}\right)\] i don't know which you are supposed to use ... but i was feeling partial to cosine

OpenStudy (paxpolaris):

the equation @myininaya had almost arrived at was also correct: \[x(t)=\frac13\sin(3t)+\cos(3t)\] but you 'd still have to do some work to arrive at Amplitude and phase.

OpenStudy (mony01):

wait how did she arrived at that?